3. Waiting times​ (in minutes) of customers in a bank where all customers enter a single waiting line and a bank where customers wait in individual lines at three different teller windows are listed below. Find the mean and median for each of the two​ samples, then compare the two sets of results.

Single Line

6.36.3

6.66.6

6.76.7

6.86.8

6.96.9

7.17.1

7.67.6

7.87.8

7.87.8

7.87.8

Individual Lines

4.14.1

5.55.5

6.06.0

6.16.1

6.26.2

7.87.8

7.87.8

8.78.7

9.29.2

10.010.0

The mean waiting time for customers in a single line is

nothing

minutes.

4. One common system for computing a grade point average​ (GPA) assigns 4 points to an​ A, 3 points to a​ B, 2 points to a​ C, 1 point to a​ D, and 0 points to an F. What is the GPA of a student who gets an A in a

33​-credit

​course, a B in each of

threethree

44​-credit

​courses, a C in a

33​-credit

​course, and a D in a

44​-credit

​course?

The mean grade point score is

nothing .

​(Round to the nearest tenth as​ needed.)

5. A​ student's course grade is based on one midterm that counts as

1515​%

of his final​ grade, one class project that counts as

2020​%

of his final​ grade, a set of homework assignments that counts as

5050​%

of his final​ grade, and a final exam that counts as

1515​%

of his final grade. His midterm score is

8282​,

his project score is

9898​,

his homework score is

7979​,

and his final exam score is

6666.

What is his overall final​ score? What letter grade did he earn​ (A, B,​ C, D, or​ F)? Assume that a mean of 90 or above is an​ A, a mean of at least 80 but less than 90 is a​ B, and so on.

His overall final score is

nothing .

​(Type an integer or a decimal rounded to one decimal place as​ needed.)

6. Five pulse rates are randomly selected from a set of measurements. The five pulse rates have a mean of

77.077.0

beats per minute. Four of the pulse rates are

8181​,

8383​,

5555​,

and

9494.

a.

Find the missing value.

b.

Suppose that you need to create a list of n values that have a specific known mean. Some of the n values can be freely selected. How many of the n values can be freely assigned before the remaining values are​ determined? (The result is referred to as the number of degrees of​ freedom.)

a. The missing value is

nothing

beats per minute.

​(Type an integer or a decimal. Do not​ round.)

7. Which of the following is always​ true?

Choose the correct answer below.

A.

Data skewed to the right have a longer left tail than right tail.

B.

For skewed​ data, the mode is farther out in the longer tail than the median.

C.

The mean and median should be used to identify the shape of the distribution.

D.

In a symmetric and​ bell-shaped distribution, the​ mean, median, and mode are the same.

8. dentify the symbols used for each of the​ following: (a) sample standard​ deviation; (b) population standard​ deviation; (c) sample​ variance; (d) population variance.

a. The symbol for sample standard deviation is

ss

s squareds2

sigmaσ

sigma squaredσ2

.

b. The symbol for population standard deviation is

ss

s squareds2

sigmaσ

sigma squaredσ2

.

c. The symbol for sample variance is

ss

s squareds2

sigmaσ

sigma squaredσ2

.

d. The symbol for population variance is

ss

s squareds2

sigmaσ

sigma squaredσ2

.

9. Listed below are the top 10 annual salaries​ (in millions of​ dollars) of TV personalities. Find the​ range, variance, and standard deviation for the sample data. Given that these are the top 10​ salaries, do we know anything about the variation of salaries of TV personalities in​ general?

4242     

4141     

4040     

3434     

2626     

2323     

2121     

1717     

16.816.8     

15.915.9

The range of the sample data is

​$nothing

million. ​(Type an integer or a​ decimal.)

10. Listed below are amounts​ (in millions of​ dollars) collected from parking meters by a security service company and other companies during similar time periods. Do the limited data listed here show evidence of stealing by the security service​ company's employees?

Security Service Company:

1.41.4

1.71.7

1.61.6

1.71.7

1.41.4

1.61.6

1.61.6

1.71.7

1.31.3

1.51.5

Open in StatCrunch

+

Copy to Clipboard

+

Open in Excel

+

Other Companies:

Find the coefficient of variation for each of the two​ samples, then compare the variation.

The coefficient of variation for the amount collected by the security service company is

nothing ​%.

​(Round to one decimal place as​ needed.)

1.91.9

1.81.8

1.71.7

1.71.7

1.61.6

1.81.8

1.61.6

1.51.5

1.91.9

1.61.6
Answer #1

3.

Converting entire data into seconds:

Single Line 23,763 25,566 26,167 26,768 27,369 26,221 29,226 30,428 30,428 30,428
Individual Lines 15,241 21,305 21,960 22,561 23,162 30,428 30,428 33,487 34,142 36,060

n1 =n2 =10

Calculation procedure: For example: 6.36.3 =6 hrs. 36 mins. 3 seconds =(6*60*60)+(36*60)+3 =21600+2160+3 =23,763 seconds.

For single line:

Mean waiting time =\bar{X1} =\Sigma X1i/n1 =27636.4 seconds

Median waiting time =M1 =27068.5 seconds

For individual lines:

Mean waiting time = \bar{X2} =\Sigma X2i/n2 =26877.4 seconds

Median waiting time =M2 =26795 seconds

Therefore, the mean and median waiting times in case of individual lines is less than that of single line.